OpenStudy (anonymous):

Solve: -2(x - 5) = -4 what is x?

OpenStudy (ash2326):

-2(x-5)=-4 -2x+10=-4 -2x=-14 x=7

jhonyy9 (jhonyy9):

-2x+10=-4 2x=14 x=14/2 =7